Thinfilms morning report 5/4/04

james haydon jhaydon at
Tue May 4 08:59:08 PDT 2004


Thinfilms Morning Report 5/4/2004



Gryphon:            ok

Innotec;            ok

Balzers :            ok

STS Dep:            ok

Implanter:            ok

Metalica:            cryo pump replacement 5/4/04-5/05/04

SCT:                completion of installation near end of May

CPD:                ok                   

WbGen-CTB:            control temp bath not working should be back up 
late next week

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            right hot pot continues to drain after drain has 
been turned off

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <>

More information about the thinfilms mailing list